# A Survey on Industrial Anomalies Synthesis

### v1

@@ 1 Introduction
(i) Limited Scope on Synthesis Methods. Existing surveys often limit their scope to specific synthesis methods, failing to provide a holistic perspective on the field. For instance,  Chen et al. emphasize data augmentation strategies but overlook the potential of more advanced synthesis methods in generating more discriminative defect representations. Similarly,  Xia et al. focus on the performance of generative adversarial networks (GAN)-based models in anomaly detction and synthesis, but fail to explore the potential of other emerging generative paradigms. This fragmented perspective restricts the understanding of how diverse methods can complement each other in addressing practical challenges.
(ii) Lack of a Dedicated Taxonomy: Existing anomaly-related surveys primarily focus on anomaly detection, often treating anomaly synthesis as a minor aspect without a structured classification. For instance,  Liu et al. broadly classify anomaly detection methods into supervised and unsupervised categories, which fails to capture the methodological diversity of anomaly synthesis. To the best of our knowledge, our survey is the first dedicated work that systematically categorizes IAS methods. By addressing this gap, we introduce a structured taxonomy tailored specifically for anomaly synthesis, providing a more comprehensive foundation for understanding and comparing different synthesis approaches.
(iii) Insufficient Emphasis on Cross-Modality Synthesis. Although existing research acknowledges the importance of leveraging multiple modalities, many review papers have not thoroughly explored how emerging technologies, such as large-scale vision-language models (VLM), are transforming the field of anomaly synthesis.  Cao et al. proposed the use of VLM for cross-modal anomaly synthesis. However, a systematic analysis of their applicability and current limitations remains urgently needed, the lack of which has hindered the exploration of how advanced methods align with the practical demands of industry.
@@
Contributions: The main contributions of this survey paper can be summarized as follows:
@@ 3 Hand-crafted Synthesis
Self-contained synthesis is a vital subset of hand-crafted synthesis. It operates by directly manipulating image regions, such as cropping, rearranging, or perturbing portions of the image, to synthesize new anomalies derived entirely from the original image. For instance,  Li et al. and  Schlüter et al. propose randomly cropping rectangular patches and pasting them back into normal images. Additionally,  Lyu et al. use Bézier curves to define the shape of anomalies, and CutPaste-based augmentation is employed to synthesize abnormal regions with the guidance of a saliency model. Moreover,  Pei et al. extract patches from the same object, apply diverse augmentations (e.g., geometric transformations, color distortions), and paste them back onto the object using a mask-guided strategy, enhancing downstream anomaly detection performance.
@@
External-dependent synthesis is an approach that synthesizes anomalies by incorporating textures from external data, such as texture libraries. Unlike self-contained methods, it overlays or blends external areas onto an otherwise anomaly-free image. For example,  Zavrtanik et al. and  Zhang et al. use Perlin noise to generate binary anomaly masks, enabling the combination of clean backgrounds with external textures to produce abnormal samples. Similarly,  Yang et al. refine the process by restricting the location of anomalies to the foreground, which further bridges the gap between synthetic and real anomalies.
@@
Inpainting-based synthesis is a specialized category within hand-crafted synthesis that synthesizes anomalies by selectively masking regions of the original image, disrupting its visual continuity [Pirnay and Chai]. Unlike methods that paste self-contained content or incorporate external textures, inpainting-based synthesis focuses on introducing missing or occluded areas, often by blacking out regions or filling them with noise. This transformation from a complete to an incomplete image is particularly effective in downstream reconstruction-based anomaly detection. For instance,  Nakanishi et al.,  Zavrtanik et al., and Li et al. synthesize abnormal samples by masking random regions and then reconstructing them using background information. In contrast,  Luo et al. develop an adaptive mask generator that selectively conceals abnormal regions while preserving the surrounding normal background, resulting in a more structured and context-aware synthesis. Additionally, Cao et al. and Tien et al. introduce random noise within the masked regions.
@@ 4 Distribution hypothesis-based Synthesis
Prior-dependent synthesis is a core approach in Distribution hypothesis-based syhthesis that leverages a geometric prior. It assumes that the latent space of normal data maps within a defined region, typically a manifold or hypersphere, while anomalies fall outside this boundary. By modeling the normal feature distribution and applying controlled perturbations, synthesized anomalies are at or beyond these borders, effectively simulating real anomaly features and improving the preformance of downstream task. For instance, Chen et al. constrain normal features within a compact space and guide synthesis via gradient ascent and truncated projection for enhanced diversity. Similarly, Chen et al. generate anomalies radially to refine decision boundaries, while Shin et al. utilize the manifold hypothesis and Gaussian annulus to perturb features and compute restoration errors for localization. In addition, Naud and Lavin embed normal data in non-Euclidean spaces and perturb along geodesic paths to synthesize anomalies.
@@
Data-driven synthesis is an effective approach within Distribution Hypothesis-Based Models that synthesizes anomalies by directly manipulating latent representations of normal data. Instead of relying on explicit prior assumptions, this method extracts latent features using models like autoencoders and then introduces controlled perturbations, such as Gaussian noise or other data-adaptive constraints, to synthesize anomalies. This flexible strategy leverages the intrinsic statistical properties of the data, producing anomalies that closely mirror distributions of abnormal features. Recent advancements explore diverse approaches to enhance the flexibility and realism of data-driven synthesis.  Liu et al. and  You et al. introduce SimpleNet and UniAD, which synthesize anomalies directly at the feature level by adding noise to extracted features to simulate abnormal features. The perturbed and original features in SimpleNet are then evaluated by a discriminator to ensure that the synthetic anomalies are distinct yet plausible. Building upon this,  Rolih et al. propose SuperSimpleNet, which confines the noise to specific regions and employs a new segmentation head, offering a more targeted anomaly synthesis strategy. Additionally,  Zavrtanik et al. develop DSR, a dual subspace re-projection network that trains a codebook and replaces the contents of masked regions in normal features by sampling from this codebook to synthesize abnormal features.
@@ 5 GM-based Synthesis
Full-image synthesis is a fundamental approach within GMs-Based Models, extensively employed for directly synthesizing industrial anomalies. It utilizes GM such as GAN and diffusion models to construct a unique mapping that transforms Gaussian noise into abnormal samples, effectively approximating the distribution of real anomalies. By training distinct models for different types of anomalies, it becomes possible to generate a wide variety of high-quality data that closely to actual anomalies.  Liu et al. establish multistage GANs that decouple texture generation and background-anomaly fusion, effectively modeling both abnormal areas and contextual coherence. To address data scarcity,  Du et al. develop Con-GAN, featuring shared data augmentation and hypersphere-based loss, it simultaneously prevent overfitting and enhance anomaly diversity.  Duan et al. also propose a two-stage GAN, namely DFMGAN, which consists of two processes: initial learning on normal samples and subsequent fine-tuning using abnormal samples. For complex anomaly synthesis,  Yang et al. innovatively combine large receptive fields for global structural modeling with patch-level refinement mechanisms, achieving synergistic integration of macro-context and micro-details.
@@
Full-image translation is a key approach within GM-Based Models, it synthesizes anomalies by transforming normal images via domain translation techniques like CycleGAN [Zhu et al.] and Pix2PixGAN [Isola et al.]. Full-image translation learns mappings between normal and anomaly domains, introducing targeted modifications (e.g. scratches, stains) while preserving structural integrity. Nowadays, it is a highly effective strategy for IAS. Recent advancements significantly improve full-image translation by enhancing the fidelity and controllability of synthesized anomalies.  Wen et al. extend CycleGAN within their anomaly detection pipeline, achieving superior performance.  Niu et al. propose SDGAN, a CycleGAN-based framework incorporating additional discriminators to refine the translation process. By focusing on modeling the distribution of abnormal samples, SDGAN enhances the realism of synthesized anomalies. Furthermore,  Zhang et al. introduce Defect-GAN, which explicitly models both the defacement and restoration processes rather than synthesizing anomalies arbitrarily. Additionally, it leverages spatial distribution maps to preserve the appearance of normal backgrounds, ensuring structural consistency in generated images.
@@
Local anomaly synthesis is a targeted approach within GMs-Based Models, designed to generate localized anomalies by selectively modifying specific regions of an image. Local anomaly synthesis typically involves removing parts of the image through annotations and employing generative models to synthesize anomaly-like textures within the masked regions. By focusing on localized transformations, it enables the creation of controlled and diverse anomaly patterns, such as scratches, holes, or surface irregularities, which closely resemble real-world anomalies. Recent advancements enhance local anomaly synthesis by aligning the mask with synthesized abnormal regions and improving spatial consistency between synthetic defects and their surrounding contexts.  Zhang et al. enhance the synthesis of realistic anomalies by amplifying their divergence from normal patterns. Their proposed RealNet perturbs the variance of denoising diffusion during reverse diffusion to synthesize a global anomaly map.  Wei et al. construct pseudo-normal backgrounds in abnormal regions to emphasize the distribution of anomalies.  Wei et al. develop DCDGAN, which trains a model on anomaly-only textures and blends them with different backgrounds, achieving multi-class and diversified anomaly synthesis.
@@ 6 VLM-based Synthesis
Single-stage synthesis leverages pre-trained VLMs with billions of parameters, leveraging their abundant feature representations and multimodal information. It synthesizes realistic anomalies and produces context-aware, highly detailed abnormal samples with minimal or no adjustments. By capitalizing on abundant cues from various modalities, single-stage synthesis balances computational efficiency and domain specificity, making it an attractive solution for industrial applications.  Sun et al. introduce a training-free anomaly synthesis method based on stable diffusion that maximizes alignment between abnormal regions and anomaly descriptions.  Hu et al. generate embeddings from diverse modalities and adjust them dynamically during the diffusion process. This strategy enables the creation of anomalies enriched with multimodal information, resulting in higher fidelity and diversity. Additionally,  Jiang et al. propose CAGEN, which employs mask annotations and text prompts to fine-tune ControlNet for precise control over the location and type of synthesized anomalies. Furthermore,  He et al. model cross-modal semantic features to allow fine-grained control over the characteristics of generated anomalies.
@@
Multi-stage synthesis is an advanced VLM technique that provides a comprehensive pipeline for anomaly synthesis. It integrates multiple processes, including anomaly synthesis and precise mask production. By iteratively refining global context and local abnormal regions, it enhances the realism, diversity, and contextual consistency of synthesized anomalies while meeting the requirements of downstream tasks better. This complete pipeline strengthens the usability and reliability of industrial systems. Recent studies contribute notably to this field.  Jin et al. propose a dual-branch VLM where one branch captures global context and the other targets abnormal regions, and it leverages a pre-trained segmentation for high-quality mask generation.  Shi et al. integrate textual guidance into global information and local synthesis, introducing a novel mask algorithm that improves small anomaly synthesis and enables precise control over intensity and direction.  Dai et al. introduce SeaS, a few-shot method that binds anomaly attributes to specific tokens using unbalanced anomaly prompts while aligning normal image features to preserve authenticity, it can also produce accurate mask annotations. Lastly,  Hu et al. develop AnomalyDiffusion, incorporating textual and positional information for precise anomaly localization and employing textual inversion to learn mask embeddings.

### v2

@@ 1 Introduction
(i) Limited Scope on Synthesis Methods. Existing surveys often limit their scope to specific synthesis methods, failing to provide a holistic perspective on the field. For instance,  chen2021surface emphasize data augmentation strategies but overlook the potential of more advanced synthesis methods in generating more discriminative defect representations. Similarly,  xia2022gan focus on the performance of generative adversarial networks (GAN)-based models in anomaly detction and synthesis, but fail to explore the potential of other emerging generative paradigms. This fragmented perspective restricts the understanding of how diverse methods can complement each other in addressing practical challenges.
(ii) Lack of a Dedicated Taxonomy: Existing anomaly-related surveys primarily focus on anomaly detection, often treating anomaly synthesis as a minor aspect without a structured classification. For instance,  liu2024deep broadly classify anomaly detection methods into supervised and unsupervised categories, which fails to capture the methodological diversity of anomaly synthesis. To the best of our knowledge, our survey is the first dedicated work that systematically categorizes IAS methods. By addressing this gap, we introduce a structured taxonomy tailored specifically for anomaly synthesis, providing a more comprehensive foundation for understanding and comparing different synthesis approaches.
(iii) Insufficient Emphasis on Cross-Modality Synthesis. Although existing research acknowledges the importance of leveraging multiple modalities, many review papers have not thoroughly explored how emerging technologies, such as large-scale vision-language models (VLM), are transforming the field of anomaly synthesis.  cao2024survey proposed the use of VLM for cross-modal anomaly synthesis. However, a systematic analysis of their applicability and current limitations remains urgently needed, the lack of which has hindered the exploration of how advanced methods align with the practical demands of industry.
@@ 3 Hand-crafted Synthesis
Self-contained synthesis is a vital subset of hand-crafted synthesis. It operates by directly manipulating image regions, such as cropping, rearranging, or perturbing portions of the image, to synthesize new anomalies derived entirely from the original image. For instance,  li2021cutpaste and  schluter2022natural propose randomly cropping rectangular patches and pasting them back into normal images. Additionally,  lyu2024reb use Bézier curves to define the shape of anomalies, and CutPaste-based augmentation is employed to synthesize abnormal regions with the guidance of a saliency model. Moreover,  pei2023self extract patches from the same object, apply diverse augmentations (e.g., geometric transformations, color distortions), and paste them back onto the object using a mask-guided strategy, enhancing downstream anomaly detection performance.
@@
External-dependent synthesis is an approach that synthesizes anomalies by incorporating textures from external data, such as texture libraries. Unlike self-contained methods, it overlays or blends external areas onto an otherwise anomaly-free image. For example,  zavrtanik2021draem and  zhang2023destseg use Perlin noise to generate binary anomaly masks, enabling the combination of clean backgrounds with external textures to produce abnormal samples. Similarly,  yang2023memseg refine the process by restricting the location of anomalies to the foreground, which further bridges the gap between synthetic and real anomalies.
@@
Inpainting-based synthesis is a specialized category within hand-crafted synthesis that synthesizes anomalies by selectively masking regions of the original image, disrupting its visual continuity [pirnay2022inpainting]. Unlike methods that paste self-contained content or incorporate external textures, inpainting-based synthesis focuses on introducing missing or occluded areas, often by blacking out regions or filling them with noise. This transformation from a complete to an incomplete image is particularly effective in downstream reconstruction-based anomaly detection. For instance,  nakanishi2021iterative,  zavrtanik2021reconstruction, and li2020superpixel synthesize abnormal samples by masking random regions and then reconstructing them using background information. In contrast,  luo2024ami develop an adaptive mask generator that selectively conceals abnormal regions while preserving the surrounding normal background, resulting in a more structured and context-aware synthesis. Additionally, cao2023collaborative and tien2023revisiting introduce random noise within the masked regions.
@@ 4 Distribution hypothesis-based Synthesis
Prior-dependent synthesis is a core approach in Distribution hypothesis-based syhthesis that leverages a geometric prior. It assumes that the latent space of normal data maps within a defined region, typically a manifold or hypersphere, while anomalies fall outside this boundary. By modeling the normal feature distribution and applying controlled perturbations, synthesized anomalies are at or beyond these borders, effectively simulating real anomaly features and improving the preformance of downstream task. For instance, chen2025unified constrain normal features within a compact space and guide synthesis via gradient ascent and truncated projection for enhanced diversity. Similarly, chen2024progressive generate anomalies radially to refine decision boundaries, while shin2023anomaly utilize the manifold hypothesis and Gaussian annulus to perturb features and compute restoration errors for localization. In addition, naud2020manifolds embed normal data in non-Euclidean spaces and perturb along geodesic paths to synthesize anomalies.
@@
Data-driven synthesis is an effective approach within Distribution Hypothesis-Based Models that synthesizes anomalies by directly manipulating latent representations of normal data. Instead of relying on explicit prior assumptions, this method extracts latent features using models like autoencoders and then introduces controlled perturbations, such as Gaussian noise or other data-adaptive constraints, to synthesize anomalies. This flexible strategy leverages the intrinsic statistical properties of the data, producing anomalies that closely mirror distributions of abnormal features. Recent advancements explore diverse approaches to enhance the flexibility and realism of data-driven synthesis.  liu2023simplenet and  you2022unified introduce SimpleNet and UniAD, which synthesize anomalies directly at the feature level by adding noise to extracted features to simulate abnormal features. The perturbed and original features in SimpleNet are then evaluated by a discriminator to ensure that the synthetic anomalies are distinct yet plausible. Building upon this,  rolih2025supersimplenet propose SuperSimpleNet, which confines the noise to specific regions and employs a new segmentation head, offering a more targeted anomaly synthesis strategy. Additionally,  zavrtanik2022dsr develop DSR, a dual subspace re-projection network that trains a codebook and replaces the contents of masked regions in normal features by sampling from this codebook to synthesize abnormal features.
@@ 5 GM-based Synthesis
Full-image synthesis is a fundamental approach within GMs-Based Models, extensively employed for directly synthesizing industrial anomalies. It utilizes GM such as GAN and diffusion models to construct a unique mapping that transforms Gaussian noise into abnormal samples, effectively approximating the distribution of real anomalies. By training distinct models for different types of anomalies, it becomes possible to generate a wide variety of high-quality data that closely to actual anomalies.  liu2019multistage establish multistage GANs that decouple texture generation and background-anomaly fusion, effectively modeling both abnormal areas and contextual coherence. To address data scarcity,  du2022new develop Con-GAN, featuring shared data augmentation and hypersphere-based loss, it simultaneously prevent overfitting and enhance anomaly diversity.  duan2023few also propose a two-stage GAN, namely DFMGAN, which consists of two processes: initial learning on normal samples and subsequent fine-tuning using abnormal samples. For complex anomaly synthesis,  yang2025defect innovatively combine large receptive fields for global structural modeling with patch-level refinement mechanisms, achieving synergistic integration of macro-context and micro-details.
@@
Full-image translation is a key approach within GM-Based Models, it synthesizes anomalies by transforming normal images via domain translation techniques like CycleGAN [CycleGAN2017] and Pix2PixGAN [pix2pix]. Full-image translation learns mappings between normal and anomaly domains, introducing targeted modifications (e.g. scratches, stains) while preserving structural integrity. Nowadays, it is a highly effective strategy for IAS. Recent advancements significantly improve full-image translation by enhancing the fidelity and controllability of synthesized anomalies.  wen2022new extend CycleGAN within their anomaly detection pipeline, achieving superior performance.  niu2020defect propose SDGAN, a CycleGAN-based framework incorporating additional discriminators to refine the translation process. By focusing on modeling the distribution of abnormal samples, SDGAN enhances the realism of synthesized anomalies. Furthermore,  zhang2021defect introduce Defect-GAN, which explicitly models both the defacement and restoration processes rather than synthesizing anomalies arbitrarily. Additionally, it leverages spatial distribution maps to preserve the appearance of normal backgrounds, ensuring structural consistency in generated images.
@@
Local anomaly synthesis is a targeted approach within GMs-Based Models, designed to generate localized anomalies by selectively modifying specific regions of an image. Local anomaly synthesis typically involves removing parts of the image through annotations and employing generative models to synthesize anomaly-like textures within the masked regions. By focusing on localized transformations, it enables the creation of controlled and diverse anomaly patterns, such as scratches, holes, or surface irregularities, which closely resemble real-world anomalies. Recent advancements enhance local anomaly synthesis by aligning the mask with synthesized abnormal regions and improving spatial consistency between synthetic defects and their surrounding contexts.  zhang2024realnet enhance the synthesis of realistic anomalies by amplifying their divergence from normal patterns. Their proposed RealNet perturbs the variance of denoising diffusion during reverse diffusion to synthesize a global anomaly map.  wei2022mask construct pseudo-normal backgrounds in abnormal regions to emphasize the distribution of anomalies.  wei2023diversified develop DCDGAN, which trains a model on anomaly-only textures and blends them with different backgrounds, achieving multi-class and diversified anomaly synthesis.
@@ 6 VLM-based Synthesis
Single-stage synthesis leverages pre-trained VLMs with billions of parameters, leveraging their abundant feature representations and multimodal information. It synthesizes realistic anomalies and produces context-aware, highly detailed abnormal samples with minimal or no adjustments. By capitalizing on abundant cues from various modalities, single-stage synthesis balances computational efficiency and domain specificity, making it an attractive solution for industrial applications.  sun2024cut introduce a training-free anomaly synthesis method based on stable diffusion that maximizes alignment between abnormal regions and anomaly descriptions.  hu2024anomalyxfusion generate embeddings from diverse modalities and adjust them dynamically during the diffusion process. This strategy enables the creation of anomalies enriched with multimodal information, resulting in higher fidelity and diversity. Additionally,  jiang2024cagen propose CAGEN, which employs mask annotations and text prompts to fine-tune ControlNet for precise control over the location and type of synthesized anomalies. Furthermore,  he2024anomalycontrol model cross-modal semantic features to allow fine-grained control over the characteristics of generated anomalies.
@@
Multi-stage synthesis is an advanced VLM technique that provides a comprehensive pipeline for anomaly synthesis. It integrates multiple processes, including anomaly synthesis and precise mask production. By iteratively refining global context and local abnormal regions, it enhances the realism, diversity, and contextual consistency of synthesized anomalies while meeting the requirements of downstream tasks better. This complete pipeline strengthens the usability and reliability of industrial systems. Recent studies contribute notably to this field.  jin2024dualanodiff propose a dual-branch VLM where one branch captures global context and the other targets abnormal regions, and it leverages a pre-trained segmentation for high-quality mask generation.  shi2025few integrate textual guidance into global information and local synthesis, introducing a novel mask algorithm that improves small anomaly synthesis and enables precise control over intensity and direction.  dai2024seas introduce SeaS, a few-shot method that binds anomaly attributes to specific tokens using unbalanced anomaly prompts while aligning normal image features to preserve authenticity, it can also produce accurate mask annotations. Lastly,  hu2024anomalydiffusion develop AnomalyDiffusion, incorporating textual and positional information for precise anomaly localization and employing textual inversion to learn mask embeddings.